1. Given a normal distribution with μ = 40 and σ = 6, find


(a) the area below 32;


(b) the area above 27;


(c) the area between 42 and 51;


(d) the x value that has 45% of the area below it;


(e) the x value that has 13% of the area above it.


1
Expert's answer
2021-07-02T11:55:07-0400

Area under a normal curve is equivalent to Probabilities.

μ = 40 and σ = 6

(a) Area below 32

Area below 32 is P(<32)

"P(<32)=P(z<\frac{32-40}{6})"

"=P(z<-1.33)"

"=0.0912"

(b) Area above 27 is P(>27)

"P(>27)=1-P(z<\frac{27-40}{6})"

"=1-P(z<-2.167)"

"=1-0.0151"

"=0.9849"

(c) Area between 42 and 51 is P(42<x<51)

"P(42<x<51)=P(<51)-P(<42)"

"=P(z<\frac{51-40}{6})-P(z<\frac{42-40}{6})"

"=P(z<1.833)-P(z<0.333)"

"=0.9666-0.6306"

"=0.336"

(d) The x value that has 45% of the area below it

"P(<x)=0.45"

The z-score corresponding to 0.45 on the left tail is -0.12566 (From z tables or =NORM.S.INV(0.45) excel function). Thus,

"\frac{X-40}{6}=-0.12566" Solving for x,

"x=6(-0.12566)+40"

"=39.246"

(e) The x value that has 13% of the area above it

"P(x>)=0.13"

"P(<x)=1-0.13=0.87"

The z-score corresponding to 0.87 on the left tail is 1.126391 (From z tables or =NORM.S.INV(0.87) excel function). Thus,

"\frac{X-40}{6}=1.126391" Solving for x,

"x=6(1.126391)+40"

"=46.758"